OpenStudy (anonymous):

the are of a rectangle is 27ft squared. and the length of the rectangle is 3 ft less than twice the width. find the dimensions of the rectangle

OpenStudy (anonymous):

L * W = 27 L = W - 3 plug the second equation in the first one: (W-3)*W = 27 W^2 - 3W - 27 = 0 solve the quadratic equation

OpenStudy (anonymous):

w=-27/2

OpenStudy (anonymous):

\[w = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a } = \frac{ -(-3) \pm \sqrt{(-3)^2 - 4(1)(-27)} }{ 2(1) } = \frac{ 3 \pm \sqrt{117} }{ 2 }\] doing this on your calculator will yield 2 answers [one for plus, one for minus] only one should make sense [one should be negative and negative length makes no sense]

OpenStudy (anonymous):

on your polynomial:\[aw^2 + bw + c = 0\]